An airplane sighted at the same time by two ground observers who are 5 miles apart and both directly west of the airplane. They report the angles of elevation as 13 degrees and 22 degrees. How high is the airplane?


1
Expert's answer
2021-07-13T13:18:23-0400

"tan 22° = (y/x) \\

y = x tan 22° = 0.404 x\\

tan 13° = {y / (x+5)}\\

0.231 = { 0.404 x / (x+5)}\\

0.231 (x+5) = 0.404 x\\

0.231 x + 1.155 = 0.404 x\\

1.155 = -0.231 x + 0.404 x\\

1.155= 0.173 x\\

x = 6.676 miles\\

Hence

Y = 0.404 x\\

Y = 0.404 (6.676)\\

Y = 2.697 miles\\"